# Attention vs Choice in Welfare Take-up: What Works for WIC?

### v3

@@ 6 Policy comparison validation
In this section, we validate our policy recommendation that the key to increasing WIC take-up is inducing choice, not raising attention among participating households. We show strong evidence from the publicly available data from Vermont WIC Program (2017), which aligns with our policy recommendations. Moreover, the pilot program was conducted on a limited budget from a WIC mini-grant, suggesting that a choice-nudging solution can be both effective and inexpensive. Both CNM and ABM interventions cost \$4000/year for the entire state of Vermont, about \$0.30 per household per year.
@@ 6.1 Site selection
The pilot program selected Barre, Brattleboro, Springfield, White River districts for CNM intervention because these four sites “reported the largest decreases in caseload in 2014, and were also the offices with the largest difference between the number of WIC-eligible children receiving Medicaid and the number of children who were enrolled in WIC” (Vermont WIC Program, 2017). [^12] Burlington was asked to test the CNM intervention because it was the “largest and most culturally diverse” site. The selection process was based on the outcome variable retention rate, implying that selected sites for the CNM treatment tend to have more downward-sloping trends in the retention rate of participating households. This coincides with our permutation test and three-period event study plot in Section 6.2. Hence, our estimate for CNM’s treatment effect is potentially conservative and understates the effectiveness of CNM. Nevertheless, we show that the treatment effect is extremely salient and statistically positive in Section 6.2.
@@
Vermont WIC Program (2017) states that it is hard to estimate the true effectiveness of the pilot program because there was a staggered rollout of EBT during the intervention period from Jun 2015 to Mar 2016. Using the EBT rollout timing from Vermont WIC Program (2015) meeting minutes, we show that the two treated groups are not different from the control groups in terms of their EBT implementation timing.
@@ 6.2 CNM treatment evaluation
(a) The red line indicates the actual $DiD_{\mathbbm{d}}$.
(b) The red line indicates the actual $PreTrend_{\mathbbm{d}}$.
@@
We plot the empirical distribution of $PreTrend_{\mathcal{D}}$ in Figure˜12. As suggested by Section˜6.1, there is a negative selection on the treated sites. There is a downward trend among the sites selected for the CNM treatment. The negative pretrend is marginally significant with a p-value of 10.1%. Hence, our treatment effect estimate will likely understate its true value due to the downward-sloping trend.
@@
In Section˜6.1, we claim that the sites are negatively selected, hence, our estimate is likely an understatement of the true treatment effect. One potential threat to the correctness of this statement is the mean reversion phenomenon. If sites that previously experienced a large decrease in retention rate are more likely to bounce back up to their mean retention rate, then the negatively selected sites will experience an increment in retention rate in the absence of the treatment. In this case, the estimated positive treatment effect does not have a causal interpretation. We address this concern by plotting Figure˜11 and taking a closer look at the retention rate changes over 2015 to 2017 for both the treated and control sites. For completeness, we make the same plot for the ABM treatment (See Figure˜11).
@@
Assuming a common mean for all 12 sites, if mean reversion is true, then the slower decline (from 2016 to 2017) in retention rate for orange lines would not necessarily indicate a positive treatment effect, but simply that those which previous decline more (from 2015-2016) bounce back to the mean of the 12 sites’ retention rates. However, if this were to be true, the lower gray lines should also exhibit similar “bounce back” behavior which we do not observe in Figure˜11. We observe the opposite, lower gray lines decline more than upper gray lines. Hence, it is unlikely that mean reversion plays a big role in our treatment effect evaluation.
@@ 6.3 ABM treatment evaluation
We test whether there is any treatment and whether there is any pretrend, both of which are two-sided hypothesis tests. We use the same methodologies to evaluate the effectiveness of the ABM treatment: permutation with a DiD estimand, canonical DiD, and event study. The results are reported in Figure˜12 and Table˜5,. All three sets of results agree that there is neither a significant pretrend nor a significant treatment effect.
Since we expect that the treatment effect of ABM kicks in immediately, we use $RR_{i,2017}-RR_{i,2016}$ instead of $RR_{i,2017}-RR_{i,2015}$. Figure˜12 shows that the actual DiD estimate is negative with a p-value $119/495\approx 24.02\%$; Figure˜12 shows that the actual pretrend estimate has a p-value $245/495\approx 49.50\%$. Our permutation tests show that both the treatment effect and the pretrend are not statistically significant.
@@
Table˜5 also show that both the pretrend and treatment effect are statistically insignificant. We plot the event study estimates and their 90% confidence intervals on Figure˜12. Both confidence intervals cover 0.
Though the treatment effect is estimated to be insignificant, weădo not claim that the ABM intervention is ineffective. Vermont WIC Program (2017) shows that ABM is useful in decreasing the rate of rescheduling and canceling appointments, decreasing workload from WIC staff. This allows the staff to focus on better serving other aspects of the program, potentially having positive long-term impacts, which are out of the scope of our study.
@@ A.1 Combining marketing and welfare take-up
We believe that this is the first work that applies the attention-choice model from the marketing literature (Hortaçsu, Madanizadeh and Puller, 2017; Heiss et al., 2021; Einav, Klopack and Mahoney, 2025) to the incomplete welfare take-up problem (see Ko and Moffitt (2024) for a recent survey on the welfare take-up literature). Ko and Moffitt (2024) lists four major mechanisms investigated by the welfare take-up literature: unawareness, limited usefulness, sign-up hassle, and stigma. We incorporate all four mechanisms into our setup: unawareness is subsumed in the attention stage, whereas the other three mechanisms are explicitly modeled in the choice stage.
@@ A.2 Inattention models
Our model introduces two innovations relative to existing two-stage inattention models in the marketing literature. First, our model does not require an exclusive shifter for the attention stage and requires only one binary exclusive shifter for the choice stage. The exclusive shifter is a binary indicator for whether a household needs to recertify its eligibility. Given that welfare programs are means-tested, such a data requirement is unlikely to be restrictive. This contrasts with the typical approach in the inattention model literature, which relies on two sets of exclusive shifters for both attention and choice stages (Heiss et al., 2021; Agarwal and Somaini, 2025). Second, we allow unobserved household-level heterogeneity to influence both attention and choice probabilities simultaneously. In contrast, most inattention models rule out unobserved heterogeneity (Honka, Hortaçsu and Vitorino, 2017; Crawford, Griffith and Iaria, 2021). We compare in detail our contribution and two leadings papers in this field.
A.2.1 Abaluck and Adams-Prassl (2021)
@@
Abaluck and Adams-Prassl (2021) exploits the Slutsky asymmetry when prices of goods vary. Importantly, the price variation of the default good (the good that inattentive consumers choose by default) has a different impact from the price variation of non-default goods. In the welfare take-up context, the default option (nonparticipation for SA) does not have a price. Hence, our context calls for a different identification strategy.
@@
We find the full model and Abaluck and Adams-Prassl (2021) similar in the constructing-with-derivative strategy for identifying functions of attention probability. Abaluck and Adams-Prassl (2021) leverages the Slutsky asymmetry in first-order derivatives (FODs) when prices vary for the default good versus for non-default good. The former affects both attention probability and choice probability; the latter affects choice probability only. Hence, their difference identifies a function of attention probability. We adopt a similar strategy to identify the ratio of attention parameters. However, looking for relevant FODs is trickier in the presence of unobserved household heterogeneity, which is ruled out in Abaluck and Adams-Prassl (2021). To accommodate household-level unobserved heterogeneity, we look for two sequences of take-up decisions that are coherent shuffle of each other. The ratio of the two relevant FODs identifies the ratio between attention parameters. The price we pay for allowing unobserved heterogeneity is the linear specification of the utility functions.
A.2.2 Agarwal and Somaini (2025)
@@
Agarwal and Somaini (2025) assumes two sets of exclusive shifters, one for the attention stage and the other for the choice stage. In contrast, our key identifying assumption is the existence of an observed fully attentive subpopulation. In our propositions, the attention shifter is not required to be exclusive but has to generate an observed fully attentive subgroup. [^13] We further modify the main identification results for the preliminary model in Appendix B to show that if the attention shifter is exclusive, exogenous and generates a fully attentive subgroup, then an exclusive choice shifter is not required. Moreover, the exclusive attention shifter in Agarwal and Somaini (2025) is assumed to have large support. In our model, the non-exclusive attention shifter is the last period binary participation status.
@@
Admittedly, though our theorems have a lower data requirement, our model setup, which requires the presence of a full attention group and a recertification process, is much more restrictive. Hence, we anticipate that readers find our work a better fit for specific empirical setups, such as welfare take-up or some other program participation contexts, whereas Agarwal and Somaini (2025)’s theoretical results apply to a wider range of topics but may impose unrealistic data assumptions in the empirical contexts that this paper specializes in.
@@ A.3 Double hurdle model
Some readers may find our model similar to the classical double hurdle model (Cragg, 1971). We clarify that our setup is substantially different from the classical double hurdle model as our outcome variable is binary, whereas Cragg (1971) explicitly states that part of the observed outcome is a “continuous and positive random variable”.
@@ A.4 Dynamic panel discrete choice with random effect
the full model introduces a dynamic panel (AR1) covariate into a generalized linear mixed-effect model (GLMM). Though such a design has been used by other works from econometrics, biometrics, and psychometrics (Wooldridge, 2005; Escaramís, Carrasco and Ascaso, 2008; Cho, Brown-Schmidt and Lee, 2018), the full model allows the AR1 term (i.e., last period take-up decision) and the random-effect component to simultaneously enter the attention and choice stages. This is a special feature of the full model that, to the best of our knowledge, has only been used by one paper: Heiss et al. (2021). However, Heiss et al. (2021) argues that the identifiability of their model comes from the presence of two sets of large-support exclusive shifters; we show that certain aspects of the model identifiability can be established semiparametrically/nonparametrically when an FA subpopulation is created by the non-exclusive attention shifter. Furthermore, the attention shifter in our paper is the binary-support lagged outcome, imposing virtually no data requirements if the data is longitudinal.
@@ C.1 Benefit imputation variable details
State groupings WIC is a federal assistance program administered by the United States Department of Agriculture (USDA). However, the actual implementation is managed at the state level. While the federal government provides a framework for the WIC program, including maximum monthly allowances for different food packages, states have some flexibility in implementing the program. For example, currently, USDA specifies that "partially breastfed infants may receive up to 104 fl. oz. of infant formula". The state government chooses which infant formula brand can be purchased with a WIC voucher/EBT card (Davis, 2012). Different infant formula brands have different prices, consequently, the dollar value of WIC benefits can vary across states within the guidelines set by the federal government.
@@
Number of kids and age of kids WIC benefits are allocated based on the nutritional needs of each child. Therefore, in general, a household with more kids will get more benefits as evidenced by Figure 13(b). Infants from birth to one year of age receive WIC food packages designed to support their growth and development during this critical period. The benefits include infant formula, infant cereal, and baby foods appropriate for their age and stage of development. Children between the ages of one and five years old receive WIC packages that focus on promoting healthy eating habits and meeting their nutritional needs as they transition to solid foods. The benefits include a variety of nutritious foods such as milk, cheese, eggs, fruits, vegetables, and whole grains. In general, the (perceived) value of food packages for infants is larger than that for children between one and five (Lora et al., 2023).
@@ D.2 Attention stage estimation results
There are two corresponding findings on the choice probability. (1) The choice probability stays between 85% to 90% during. The constant choice probability of households from each education level is determined by the model design. The choice probability surges to 95% when the baby is delivered, and then it steadily declines as LeftBF decreases to 0. (2) Higher-educated households have higher choice probability. This finding aligns with some empirical research that higher-educated parents value nutrition for their children more than lower-educated parents Bere et al. (2008). Next, we show that the full model outputs the same qualitative results.
@@ Appendix E Additional findings from the full model
E.1 Targeting pattern in comparison to Finkelstein and Notowidigdo (2019)
Finkelstein and Notowidigdo (2019) finds that intervention that aims at promoting welfare take-up tend to help less needy households who receive lower benefits from the program and are likely to be less socioeconomically disadvantaged. We simulate the policy intervention in Finkelstein and Notowidigdo (2019) as a one-shot attention boost by 30% for all households at period 1.
We summarize in Table˜7 the average characteristics of all participating households under different interventions. The average observed characteristics, such as benefit, number of kids below one, education, and local accessibility, virtually do not vary across these four cases. In other words, we do not find the same targeting pattern in the households’ observed characteristics. Instead, we find a targeting pattern on the random effect when the policy increases the take-up rate substantially. We provide a unifying explanation for the targeting on observable patterns in Finkelstein and Notowidigdo (2019) and the targeting on unobservable patterns in our paper.
@@ E.1 Targeting pattern in comparison to Finkelstein and Notowidigdo (2019)
Both types of targeting tend to promote take-up more among households with lower choice probability. In Finkelstein and Notowidigdo (2019), households with lower benefits and lower needs for welfare benefits have a lower choice tendency. On the other hand, a lower random effect also leads to a lower choice tendency. Hence, we can unify the two targeting patterns as helping out households with lower choice tendency. This unifying explanation is intuitive. Households with higher choice tendency are more likely to join the program in the absence of any intervention. Any intervention is likely to induce more participation among households with lower choice tendency.

### v5

@@ 6 Policy comparison validation
In this section, we validate our policy recommendation that the key to increasing WIC take-up is inducing choice, not raising attention among participating households. We show strong evidence from the publicly available data from vermont2017wic2five, which aligns with our policy recommendations. Moreover, the pilot program was conducted on a limited budget from a WIC mini-grant, suggesting that a choice-nudging solution can be both effective and inexpensive. Both CNM and ABM interventions cost \$4000/year for the entire state of Vermont, about \$0.30 per household per year.
@@ 6.1 Site selection
The pilot program selected Barre, Brattleboro, Springfield, White River districts for CNM intervention because these four sites “reported the largest decreases in caseload in 2014, and were also the offices with the largest difference between the number of WIC-eligible children receiving Medicaid and the number of children who were enrolled in WIC” (vermont2017wic2five). [^10] Burlington was asked to test the CNM intervention because it was the “largest and most culturally diverse” site. The selection process was based on the outcome variable retention rate, implying that selected sites for the CNM treatment tend to have more downward-sloping trends in the retention rate of participating households. This coincides with our permutation test and three-period event study plot in Section 6.2. Hence, our estimate for CNM’s treatment effect is potentially conservative and understates the effectiveness of CNM. Nevertheless, we show that the treatment effect is extremely salient and statistically positive in Section 6.2.
@@
vermont2017wic2five states that it is hard to estimate the true effectiveness of the pilot program because there was a staggered rollout of EBT during the intervention period from Jun 2015 to Mar 2016. Using the EBT rollout timing from vermontwic2015minutes meeting minutes, we show that the two treated groups are not different from the control groups in terms of their EBT implementation timing.
@@ 6.2 CNM treatment evaluation
(a) The black line indicates the actual $DiD_{\mathbbm{d}}$.
(b) The black line indicates the actual $PreTrend_{\mathbbm{d}}$.
@@
We plot the empirical distribution of $PreTrend_{\mathcal{D}}$ in Figure˜14. As suggested by Section˜6.1, there is a negative selection on the treated sites. There is a downward trend among the sites selected for the CNM treatment. The negative pretrend is marginally significant with a p-value of 10.1%. Hence, our treatment effect estimate will likely understate its true value due to the downward-sloping trend.
@@
In Section˜6.1, we claim that the sites are negatively selected, hence, our estimate is likely an understatement of the true treatment effect. One potential threat to the correctness of this statement is the mean reversion phenomenon. If sites that previously experienced a large decrease in retention rate are more likely to bounce back up to their mean retention rate, then the negatively selected sites will experience an increment in retention rate in the absence of the treatment. In this case, the estimated positive treatment effect does not have a causal interpretation. We address this concern by plotting Figure˜13 and taking a closer look at the retention rate changes over 2015 to 2017 for both the treated and control sites. For completeness, we make the same plot for the ABM treatment (See Figure˜13).
@@
Assuming a common mean for all 12 sites, if mean reversion is true, then the slower decline (from 2016 to 2017) in retention rate for orange lines would not necessarily indicate a positive treatment effect, but simply that those which previous decline more (from 2015-2016) bounce back to the mean of the 12 sites’ retention rates. However, if this were to be true, the lower gray lines should also exhibit similar “bounce back” behavior which we do not observe in Figure˜13. We observe the opposite, lower gray lines decline more than upper gray lines. Hence, it is unlikely that mean reversion plays a big role in our treatment effect evaluation.
@@ 6.3 ABM treatment evaluation
We test whether there is any treatment and whether there is any pretrend, both of which are two-sided hypothesis tests. We use the same methodologies to evaluate the effectiveness of the ABM treatment: permutation with a DiD estimand, canonical DiD, and event study. The results are reported in Figure˜14 and Table˜5,. All three sets of results agree that there is neither a significant pretrend nor a significant treatment effect.
Since we expect that the treatment effect of ABM kicks in immediately, we use $RR_{i,2017}-RR_{i,2016}$ instead of $RR_{i,2017}-RR_{i,2015}$. Figure˜14 shows that the actual DiD estimate is negative with a p-value $119/495\approx 24.02\%$; Figure˜14 shows that the actual pretrend estimate has a p-value $245/495\approx 49.50\%$. Our permutation tests show that both the treatment effect and the pretrend are not statistically significant.
@@
Table˜5 also show that both the pretrend and treatment effect are statistically insignificant. We plot the event study estimates and their 90% confidence intervals on Figure˜14. Both confidence intervals cover 0.
Though the treatment effect is estimated to be insignificant, weădo not claim that the ABM intervention is ineffective. vermont2017wic2five shows that ABM is useful in decreasing the rate of rescheduling and canceling appointments, decreasing workload from WIC staff. This allows the staff to focus on better serving other aspects of the program, potentially having positive long-term impacts, which are out of the scope of our study.
@@ A.1 Combining marketing and welfare take-up
We believe that this is the first work that applies the attention-choice model from the marketing literature (hortaccsu2017power; heiss2021inattention; einav2025selling) to the incomplete welfare take-up problem (see ko2024take for a recent survey on the welfare take-up literature). ko2024take lists four major mechanisms investigated by the welfare take-up literature: unawareness, limited usefulness, sign-up hassle, and stigma. We incorporate all four mechanisms into our setup: unawareness is subsumed in the attention stage, whereas the other three mechanisms are explicitly modeled in the choice stage.
@@ A.2 Inattention models
Our model introduces two innovations relative to existing two-stage inattention models in the marketing literature. First, our model does not require an exclusive shifter for the attention stage and requires only one binary exclusive shifter for the choice stage. The exclusive shifter is a binary indicator for whether a household needs to recertify its eligibility. Given that welfare programs are means-tested, such a data requirement is unlikely to be restrictive. This contrasts with the typical approach in the inattention model literature, which relies on two sets of exclusive shifters for both attention and choice stages (heiss2021inattention; agarwal2025demand). Second, we allow unobserved household-level heterogeneity to influence both attention and choice probabilities simultaneously. In contrast, most inattention models rule out unobserved heterogeneity (honka2017advertising; crawford2021survey). We compare in detail our contribution and two leadings papers in this field.
A.2.1 abaluck2021consumers
@@
abaluck2021consumers exploits the Slutsky asymmetry when prices of goods vary. Importantly, the price variation of the default good (the good that inattentive consumers choose by default) has a different impact from the price variation of non-default goods. In the welfare take-up context, the default option (nonparticipation for SA) does not have a price. Hence, our context calls for a different identification strategy.
@@
We find the full model and abaluck2021consumers similar in the constructing-with-derivative strategy for identifying functions of attention probability. abaluck2021consumers leverages the Slutsky asymmetry in first-order derivatives (FODs) when prices vary for the default good versus for non-default good. The former affects both attention probability and choice probability; the latter affects choice probability only. Hence, their difference identifies a function of attention probability. We adopt a similar strategy to identify the ratio of attention parameters. However, looking for relevant FODs is trickier in the presence of unobserved household heterogeneity, which is ruled out in abaluck2021consumers. To accommodate household-level unobserved heterogeneity, we look for two sequences of take-up decisions that are coherent shuffle of each other. The ratio of the two relevant FODs identifies the ratio between attention parameters. The price we pay for allowing unobserved heterogeneity is the linear specification of the utility functions.
A.2.2 agarwal2025demand
@@
agarwal2025demand assumes two sets of exclusive shifters, one for the attention stage and the other for the choice stage. In contrast, our key identifying assumption is the existence of an observed fully attentive subpopulation. In our propositions, the attention shifter is not required to be exclusive but has to generate an observed fully attentive subgroup. [^11] We further modify the main identification results for the preliminary model in Appendix B to show that if the attention shifter is exclusive, exogenous and generates a fully attentive subgroup, then an exclusive choice shifter is not required. Moreover, the exclusive attention shifter in agarwal2025demand is assumed to have large support. In our model, the non-exclusive attention shifter is the last period binary participation status.
@@
Admittedly, though our theorems have a lower data requirement, our model setup, which requires the presence of a full attention group and a recertification process, is much more restrictive. Hence, we anticipate that readers find our work a better fit for specific empirical setups, such as welfare take-up or some other program participation contexts, whereas agarwal2025demand’s theoretical results apply to a wider range of topics but may impose unrealistic data assumptions in the empirical contexts that this paper specializes in.
@@ A.3 Double hurdle model
Some readers may find our model similar to the classical double hurdle model (cragg1971some). We clarify that our setup is substantially different from the classical double hurdle model as our outcome variable is binary, whereas cragg1971some explicitly states that part of the observed outcome is a “continuous and positive random variable”.
@@ A.4 Dynamic panel discrete choice with random effect
The full model introduces a dynamic panel (AR1) covariate into a generalized linear mixed-effect model (GLMM). Though such a design has been used by other works from econometrics, biometrics, and psychometrics (wooldridge2005simple; escaramis2008detection; cho2018autoregressive), the full model allows the AR1 term (i.e., last period take-up decision) and the random-effect component to simultaneously enter the attention and choice stages. This is a special feature of the full model that, to the best of our knowledge, has only been used by one paper: heiss2021inattention. However, heiss2021inattention argues that the identifiability of their model comes from the presence of two sets of large-support exclusive shifters; we show that certain aspects of the model identifiability can be established semiparametrically/nonparametrically when an FA subpopulation is created by the non-exclusive attention shifter. Furthermore, the attention shifter in our paper is the binary-support lagged outcome, imposing virtually no data requirements if the data is longitudinal.
@@ C.1 Benefit imputation variable details
State groupings WIC is a federal assistance program administered by the United States Department of Agriculture (USDA). However, the actual implementation is managed at the state level. While the federal government provides a framework for the WIC program, including maximum monthly allowances for different food packages, states have some flexibility in implementing the program. For example, currently, USDA specifies that "partially breastfed infants may receive up to 104 fl. oz. of infant formula". The state government chooses which infant formula brand can be purchased with a WIC voucher/EBT card (davis2012bidding). Different infant formula brands have different prices, consequently, the dollar value of WIC benefits can vary across states within the guidelines set by the federal government.
@@
Number of kids and age of kids WIC benefits are allocated based on the nutritional needs of each child. Therefore, in general, a household with more kids will get more benefits as evidenced by Figure 15(b). Infants from birth to one year of age receive WIC food packages designed to support their growth and development during this critical period. The benefits include infant formula, infant cereal, and baby foods appropriate for their age and stage of development. Children between the ages of one and five years old receive WIC packages that focus on promoting healthy eating habits and meeting their nutritional needs as they transition to solid foods. The benefits include a variety of nutritious foods such as milk, cheese, eggs, fruits, vegetables, and whole grains. In general, the (perceived) value of food packages for infants is larger than that for children between one and five (lora2023factors).
@@ D.2 Attention stage estimation results
There are two corresponding findings on the choice probability. (1) The choice probability stays between 85% to 90% during. The constant choice probability of households from each education level is determined by the model design. The choice probability surges to 95% when the baby is delivered, and then it steadily declines as LeftBF decreases to 0. (2) Higher-educated households have higher choice probability. This finding aligns with some empirical research that higher-educated parents value nutrition for their children more than lower-educated parents bere2008parents. Next, we show that the full model outputs the same qualitative results.
@@ Appendix E Additional findings from the full model
E.1 Targeting pattern in comparison to finkelstein2019take
finkelstein2019take finds that intervention that aims at promoting welfare take-up tend to help less needy households who receive lower benefits from the program and are likely to be less socioeconomically disadvantaged. We simulate the policy intervention in finkelstein2019take as a one-shot attention boost by 30% for all households at period 1.
We summarize in Table˜7 the average characteristics of all participating households under different interventions. The average observed characteristics, such as benefit, number of kids below one, education, and local accessibility, virtually do not vary across these four cases. In other words, we do not find the same targeting pattern in the households’ observed characteristics. Instead, we find a targeting pattern on the random effect when the policy increases the take-up rate substantially. We provide a unifying explanation for the targeting on observable patterns in finkelstein2019take and the targeting on unobservable patterns in our paper.
@@ E.1 Targeting pattern in comparison to finkelstein2019take
Both types of targeting tend to promote take-up more among households with lower choice probability. In finkelstein2019take, households with lower benefits and lower needs for welfare benefits have a lower choice tendency. On the other hand, a lower random effect also leads to a lower choice tendency. Hence, we can unify the two targeting patterns as helping out households with lower choice tendency. This unifying explanation is intuitive. Households with higher choice tendency are more likely to join the program in the absence of any intervention. Any intervention is likely to induce more participation among households with lower choice tendency.